OpenStudy (anonymous):

Which rational number is one-half of the way between 2 and 6.5?

zepdrix (zepdrix):

Do you know how to find the distance between 2 and 6.5? :)

OpenStudy (anonymous):

4.75 2.25 4.25 4.5

OpenStudy (anonymous):

4.5 would be the answer to me?

zepdrix (zepdrix):

4.5 is the space between 2 and 6.5, the full distance :) good. They want half of that.

OpenStudy (anonymous):

2.25

zepdrix (zepdrix):

yay good job \c:/

zepdrix (zepdrix):

oh oh oh wait wait im being stupid :O

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

and I thought "Of wow thats simple.." lol

zepdrix (zepdrix):

So you found half of the distance between 2 and 6.5. Good 2.25. We started at 2 though. So you need to add 2.25 to your 2 to find the half-way point.

OpenStudy (anonymous):

oh*

OpenStudy (anonymous):

4.25

zepdrix (zepdrix):

2.25 is clearly not the half-way point lol

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

yeah lol!

OpenStudy (anonymous):

so its 4.25?

jimthompson5910 (jim_thompson5910):

An alternative is to add up the two numbers. Then divide by 2 \[\Large \frac{2+6.5}{2}=\frac{8.5}{2} = 4.25\] and you get the same answer. The method zepdrix used is probably more intuitive though.

zepdrix (zepdrix):

ahh average :) ya i like that better :D

OpenStudy (anonymous):

Thanks guys!:)